# Online Housing Market

### v1

@@ 7.2 Partition Functions Based on Schedules
Consider one last time instance $I$ described in the left part of Figure 4. Let me compute the earliest departure partition $\zeta(I)$ by running the modified version of Algorithm 8 described in the previous paragraph. At $d_{1}$, during the earliest departure time of an agent, $N_{<d_{1}}$ contains agents 1, 2, and 3. Therefore, $\{1\}$ and $\{2,3\}$ are added to the partition. All the other agents will remain alone in the partition. Thus, the algorithm halts and returns the earliest departure partition $\zeta(I)=\{\{1\},\{2,3\},\{4\},\{5\}\}$, which is identical to the departing agent excluded partition $\gamma(I)$. As a consequence, the matching returned by Algorithm 6 using the earliest departure partition $\zeta$ is the matching $M$ depicted in Figure 4.
@@
Algorithm 6 using the earliest departure partition $\zeta$ is $SIC$.
@@
Note first that, since the earliest departure partition $\zeta$ is $SOC$, I know by Propositions 16 that Algorithm 6 using it as input is $WIC$. It remains to show that no agent has an incentive to misreport either her arrival or departure times. Let $S^{\prime}$ denote the only coalition of the earliest departure partition $\zeta$ that may contain more than one member. Note that, according to the modification performed to line 6 of Algorithm 8 and the input scheduling $\xi$ such that $\xi_{1}=T$, $S^{\prime}=S\setminus\{\delta(1)\}$, where $S=N_{<\delta(1)}$. I know that for any agent $j$ that does not belong to $S^{\prime}$, either (i) $j=\delta(1)$ or (ii) $a_{j}>d_{\delta(1)}$ hold. In both cases, agent $j$ cannot misreport her departure time $d_{j}$ to $d^{\prime}_{j}$ in order to change this fact since $d^{\prime}_{j}<d_{j}$ (for case (i)), and she cannot misreport her arrival time $a_{j}$ to $a^{\prime}_{j}$ since $a^{\prime}_{j}>a_{j}$ (for case (ii)). Therefore, these agents have no incentive to misreport since they will keep their items in any case during Algorithm 6. Consider now the case of an agent $j$ belonging to $S^{\prime}$. The only way for agent $j$ to change the outcome of the earliest departure partition $\zeta$ is either to declare a departure time $d^{\prime}_{j}$ such that $d^{\prime}_{j}<d_{\delta(1)}$, or an arrival time $a^{\prime}_{j}$ such that $a^{\prime}_{j}>d_{\delta(1)}$. In both cases, the result for agent $j$ is that she will be left alone by the earliest departure partition $\zeta$, and she will keep her item during Algorithm 6. Since TTC is $IR$, agent $j$ has no incentive to misreport since she should receive an item that is at least as preferred as her own. Therefore, no agent has an incentive to misreport either her arrival or departure time, and the procedure is $SIC$. ∎
The main drawback of using the earliest departure partition $\zeta$ as input of Algorithm 6 is that only one coalition of $\zeta$ contains more than one agent, and the exchanges performed during the TTC procedure may involve very few agents. The number of agents involved in the TTC procedure highly depends on the set of agents present in the market at the time when the first agent leaves, and its range is from 1 (when $a_{2}>d_{1}$) to $n-1$ (when $a_{n}<d_{\delta(1)}$).
@@ 8 Conclusion and Future Works
I extended the serial dictatorship and top trading cycle procedures to an online setting, aiming to develop mechanisms that are Pareto-efficient, individually rational, and incentivize agents to truthfully reveal their preferences, as well as their actual arrival and departure times. Several variants of these mechanisms were proposed and are summarized in Table 1 along with their respective properties. It is important to note that Table 1 alone may not suffice to fully compare different mechanisms. For instance, the static and dynamic versions of the serial dictatorship procedure using the ascending arrival permutation $\alpha$ (corresponding to rows 2 and 3 in Table 1) appear incomparable based solely on their listed properties. However, the allocation returned by the dynamic version is never Pareto-dominated by the allocation produced by the static version, while the reverse may hold, as demonstrated in Example 4. This observation suggests that the dynamic version is more efficient in practice than the static version. Another example involves the static top trading cycle procedure (Algorithm 6) using the earliest departure partition $\zeta$, which seems to dominate other partition functions since it satisfies a broader set of properties. However, $\zeta$ typically identifies only a single subset of agents where exchanges can occur. The size of this subset is highly dependent on the departure time of the earliest agent to leave, making the mechanism less predictable and potentially inefficient in markets where agents frequently arrive and depart.
@@
Algorithm 6 with earliest $\surd$ $\surd$ $\surd$ $\surd$
departure partition $\zeta$ (Prop. 16) (Prop. 23) (Prop. 23)
The paper also presents additional results not summarized in Table 1. For example, I proved that the only mechanism guaranteeing a $\mathcal{M}(I)$-Pareto-optimal ($\mathcal{S}(I)$-Pareto-optimal, respectively) allocation for any instance $I$ is the static serial dictatorship procedure (safe serial dictatorship procedure, respectively) using the ascending departure permutation $\delta$ as input (Propositions 4 and 14, respectively). As a consequence, $\mathcal{M}$-$PO$ is incompatible with $IR$, $a$-$IC$, and $d$-$IC$ (Corollary 13). Additionally, I demonstrated that the serial dictatorship procedure is inherently incompatible with $a$-$IC$, regardless of the input permutation function (Propositions 8 and 12). Moreover, I established that the ascending departure permutation is the only input permutation function that ensures the serial dictatorship procedure satisfies $d$-$IC$ (Propositions 7 and 11). Finally, I showed that no permutation function can make the safe serial dictatorship procedure $WIC$ (Proposition 15).
An intriguing extension of this work would be to develop dynamic versions of the TTC procedure, allowing agents to participate multiple times during the execution of Algorithm 6, each time with their currently allocated item. Another promising avenue is to explore the design of strongly incentive-compatible mechanisms that facilitate more exchanges than Algorithm 6 under the earliest departure partition $\zeta$. Additionally, relaxing my assumptions about the design of permutation and partition functions—specifically allowing them to depend on parameters beyond just the arrival and departure times of agents—opens another interesting research direction. Finally, investigating relaxed forms of strategy-proofness could provide valuable insights, particularly by accounting for agents’ limited foresight regarding future arrivals of items and agents. Such an approach could incorporate probabilistic assumptions about arrival patterns, enabling mechanisms to achieve improved allocations. If the decision-maker has knowledge of these probabilities, optimal decisions could be crafted based on this information. These types of probabilistic assumptions are closely related to the online stochastic matching problem (?, ?). A similar approach has already been considered for the fair assignment of public goods (?, ?).

### v2

@@ 7.3 The Threshold-Triggered Partition
Consider one last time instance $I$ described in the left part of Figure 4. Let us compute the threshold-triggered partition $\zeta(I)$ by running Algorithm 9 with threshold $\tau=t^{-}$. Note that this means the first agent to leave after the threshold will be the first to exit the market. At $d_{1}$, during the earliest departure time of an agent, $N_{<d_{1}}$ contains agents 1, 2, and 3. Therefore, $\{1\}$ and $\{2,3\}$ are added to the partition. All the other agents will remain alone in the partition. Thus, the algorithm halts and returns the threshold-triggered partition $\zeta(I)=\{\{1\},\{2,3\},\{4\},\{5\}\}$, which is identical to the departing agent excluded partition $\gamma(I)$. As a consequence, the matching returned by Algorithm 6 using the threshold-triggered partition $\zeta$ is the matching $M$ depicted in Figure 4.
Note that the outcome of the algorithm may vary depending on the threshold used. For example, consider a new threshold $\tau^{\prime}$ such that $a_{4}<\tau^{\prime}<d_{2}$. In this case, the first agent to leave after the threshold is agent 2 at time $d_{2}$. The resulting partition is then $\zeta(I)=\{\{1\},\{2\},\{3,4\},\{5\}\}$. This partition is almost identical to the scheduled departure partition of Example 7, except that agents 1 and 2 do not belong to the same subset. However, since these two agents did not exchange their houses during the execution of Algorithm 6, the outcome returned by Algorithm 6 for this partition $\zeta(I)$ remains $M^{\prime}$.
@@
Algorithm 6 using the threshold-triggered partition $\zeta$ is $SIC$.
@@
Note first that, since the threshold-triggered partition $\zeta$ is $PP$, we know from Proposition 16 that Algorithm 6, when using $\zeta$ as input, is $WIC$. It remains to show that no agent has an incentive to misreport either her arrival or departure time. Let $S$ denote the only coalition in the threshold-triggered partition $\zeta$ that may contain more than one agent, and let $k$ be the agent whose departure time is the earliest among those occurring after the threshold $\tau$. According to Algorithm 9, we have $S=N_{<d_{k}}\setminus(B\cup\{k\})$, where $B$ is the set of agents whose departure times are earlier than $\tau$. For any agent $j\notin S$, one of the following must hold:
  1. (i)
    $j=k$;
  2. (ii)
    $a_{j}>d_{k}$;
  3. (iii)
    $d_{j}<\tau$.
In cases (i) and (iii), agent $j$ cannot alter her partition membership by misreporting her departure time $d_{j}$ as $d^{\prime}_{j}$, since this would require $d^{\prime}_{j}<d_{j}$. Indeed, this would imply $d^{\prime}_{j}<d_{j}<\tau$ in case (iii); and in case (i), either $d^{\prime}_{k}<\tau$, in which case agent $k$ would leave before the threshold and be excluded from the coalition, or $d^{\prime}_{k}\geq\tau$ and $d^{\prime}_{k}$ would remain the earliest departure after $\tau$, so the partition would remain unchanged. In case (ii), agent $j$ cannot manipulate her arrival time to join the coalition, since $a^{\prime}_{j}>a_{j}$ would imply $a^{\prime}_{j}>d_{k}$, and therefore her arrival would occur after agent $k$ has departed. Finally, misreporting her arrival time in cases (i) and (iii), or her departure time in case (ii), has no impact on her coalition, which remains a singleton. Thus, agents not in $S$ have no incentive to misreport: they keep their own item under Algorithm 6, regardless of their report.
Now consider an agent $j\in S$. The only way for her to alter the partition formed by the threshold-triggered partition $\zeta$ is to declare a departure time $d^{\prime}_{j}<d_{k}$ (either $d^{\prime}_{j}<\tau$, or $d^{\prime}_{j}\in[\tau,d_{k})$), or an arrival time $a^{\prime}_{j}>d_{k}$. In both cases, agent $j$ would be placed in a singleton set in the partition and would therefore keep her own item under Algorithm 6. Since TTC is individually rational ($IR$), agent $j$ has no incentive to misreport: she would receive an item at least as preferred as her own under truthful reporting. Hence, no agent has an incentive to misreport either her arrival or departure time, and the procedure is $SIC$. ∎
The main drawback of using the threshold-triggered partition $\zeta$ as input to Algorithm 6 is that only one coalition in $\zeta$ may contain more than one agent, and the exchanges performed during the TTC procedure may therefore involve very few agents. The number of agents participating in the TTC procedure depends heavily on the set of agents present in the market at the earliest departure time occurring after the threshold, and this number can range from $1$ to $n-1$.
@@ 8 Conclusion and Future Works
We extended the serial dictatorship and top trading cycle procedures to an online setting, aiming to develop mechanisms that are Pareto-efficient, individually rational, and incentivize agents to truthfully reveal their preferences, as well as their actual arrival and departure times. Several variants of these mechanisms were proposed and are summarized in Table 1 along with their respective properties. It is important to note that Table 1 alone may not suffice to fully compare different mechanisms. For instance, the static and dynamic versions of the serial dictatorship procedure using the ascending arrival permutation $\alpha$ (corresponding to rows 2 and 3 in Table 1) appear incomparable based solely on their listed properties. However, the allocation returned by the dynamic version is never Pareto-dominated by the allocation produced by the static version, while the reverse may hold, as demonstrated in Example 4. This observation suggests that the dynamic version is more efficient in practice than the static version. Another example involves the static top trading cycle procedure (Algorithm 6) using the threshold-triggered partition $\zeta$, which seems to dominate other partition functions since it satisfies a broader set of properties. However, $\zeta$ typically identifies only a single subset of agents where exchanges can occur. The size of this subset is highly dependent on the departure time of the earliest agent to leave after the threshold, making the mechanism less predictable and potentially inefficient in markets where agents frequently arrive and depart.
@@
Algorithm 6 with threshold- $\surd$ $\surd$ $\surd$ $\surd$
triggered partition $\zeta$ (Prop. 16) (Prop. 23) (Prop. 23)
The paper also presents additional results not summarized in Table 1. For example, we proved that the only mechanism guaranteeing a $\mathcal{M}(I)$-Pareto-optimal ($\mathcal{S}(I)$-Pareto-optimal, respectively) allocation for any instance $I$ is the static serial dictatorship procedure (safe serial dictatorship procedure, respectively) using the ascending departure permutation $\delta$ as input (Propositions 4 and 14, respectively). As a consequence, $\mathcal{M}(I)$-$PO$ is incompatible with $IR$, $a$-$IC$, and $d$-$IC$ (Corollary 13). Additionally, we demonstrated that the serial dictatorship procedure is inherently incompatible with $a$-$IC$, regardless of the input permutation function (Propositions 8 and 12). Moreover, we established that the ascending departure permutation is the only input permutation function that ensures the serial dictatorship procedure satisfies $d$-$IC$ (Propositions 7 and 11). Finally, we showed that no permutation function can make the safe serial dictatorship procedure $WIC$ (Proposition 15).
An intriguing extension of this work would be to develop dynamic versions of the TTC procedure, allowing agents to participate multiple times during the execution of Algorithm 6, each time with their currently allocated item. Another promising avenue is to explore the design of strongly incentive-compatible mechanisms that facilitate more exchanges than Algorithm 6 under the threshold-triggered partition $\zeta$. Additionally, relaxing our assumptions about the design of permutation and partition functions—specifically allowing them to depend on parameters beyond just the arrival and departure times of agents—opens another interesting research direction. Finally, investigating relaxed forms of strategy-proofness could provide valuable insights, particularly by accounting for agents’ limited foresight regarding future arrivals of items and agents. Such an approach could incorporate probabilistic assumptions about arrival patterns, enabling mechanisms to achieve improved allocations. If the decision-maker has knowledge of these probabilities, optimal decisions could be crafted based on this information. These types of probabilistic assumptions are closely related to the online stochastic matching problem (?, ?). A similar approach has already been considered for the fair assignment of public goods (?, ?).